Find the density of air at 13.79 MPaa and 172.8 K, using (a) pv = RT, (5) pv = ZRT, (c) van der Waals equation. (d) Which of the three answers do you believe is more accurate?

1
Expert's answer
2021-12-13T11:10:10-0500

PV= RT

"V=\frac{RT}{P}"

"V=\frac{8.31\times172.8}{13.79\times10^6}"

"V=1.041\times10^{-4}m^3"

Now

"PV=RT"



"13.79\times1.041\times10^{-4}={8.31\times172.8}"

LHS=RHS

(2)

"Z=\frac{PV}{RT}"

"Z=\frac{13.79\times10^6\times1.041\times10^{-4}}{8.31\times172.8}=0.99"

PV=ZRT

LHS=RHS

(3)

Vander Wal equation

"(P-\frac{an^2}{V^2})(V-nb)=nRT"


"(13.79\times10^6-\frac{a}{1.014\times10^{-4}})(1.014\times10^{-4}-b)=1\times8.321 \times172.8"

LHS=RHS